You uncle heard that real estate investment could yield remarkable return for investor, and he told you that he is offered a mortgage loan of RM 450, 000 with a rate of 4.5% by Bank A to purchase a double-storey terrace house in Semenyih, the loan is scheduled to be repaid in 35 years with a fixed monthly payment. Use the knowledge that you learned in Finance class,


a.               Calculate the monthly fixed payment to the bank.


b.              Use an amortization table to explain to your uncle the total interest rate that he is going to pay to the bank if he decides to accept the loan.


1
Expert's answer
2022-01-11T11:19:27-0500

a.

"A=S(i+\frac{i}{(1+i)^n-1})"

S=450 000

"i=\frac{0.045}{12}=0.00375"

"n=35\times12=420"

"A=S(i+\frac{i}{(1+i)^n-1})=450 000(0.00375+\frac{0.00375}{(1.00375)^{420}-1}=2129.66"


b.

"i=\frac{2129.66\times420}{450 000}=1.987678" 198.767%